# A New Metric Function for SC-based Polar Decoders: Polarization, Pruning, and Fast Decoders

### v2

@@ IV Metric Polarization Estimation
In this section, we numerically evaluate the path metrics of FSCL decoding for PAC codes across different rate profiles. For a PAC$(64,32)$ code operating at $E_{b}/N_{0}=2.5~{}$dB, $\mathbb{E}[\phi(X;Y)]=0.7944$. For the Monte Carlo (MC)-based rate profile obtained in [32], Fig. 9 shows the special nodes corresponding to this code. A conventional SCL decoder would need to visit $1262N-2=$ 126 nodes of the polarization tree. However, FSCL significantly reduces this requirement to only $22$ node visits, resulting in more than a 5-fold decrease in the number of node visits. In this figure, we also show the corresponding average value of the polarized metric functions obtained from the polarization of the mutual information $I(W)$ at each node. The leftmost leaf node of the tree is a Rate-0 node and contains 8 LLR values, leading to 8 metric functions, all of which have an average value of $0.1891$. On the other hand, the rightmost leaf node shows an average metric function value of $1$, indicating that these are reliable channels, and the decoder is most likely to detect the correct branch for them.
@@
Figure 10: Comparison of the partial optimum average path metric corresponding to the correctly decoded codewords of fast SC decoding of PAC$(64,32)$ code with the corresponding partially polarized channel capacity rate profile over BI-AWGN at $E_{b}/N_{0}=2.5~{}$dB with MC rate profile.
@@
Fig. 10 shows the sample average path metric values for the FSCL decoding of the PAC$(64,32)$ code, based on $10^{4}$ samples of correctly decoded codewords. This figure also includes the cumulative values of the average of our proposed polarized metric values derived from the polarization tree shown in Fig. 9. It can be observed that the proposed partially polarized average metric provides an excellent estimation of the actual sample average metric function values. This suggests that these values can be effectively used for normalizing metric function values in various practical settings as well.
@@
Figure 11: Correct path metric sample variance corresponding to the correctly decoded codewords of fast SC decoding of PAC$(64,32)$ code of Fig. 10 over BI-AWGN at $E_{b}/N_{0}=2.5~{}$dB with MC rate profile.
@@
Fig. 11 also demonstrates the metric variance of the PAC$(64,32)$ from Fig. 10 for the data bits. We observe that the variance at the data bits is nearly zero, suggesting that the average value of the metric function serves as an effective approximation.
@@
Figure 12: Comparison of the partial optimum average path metric corresponding to the correctly decoded codewords of fast SC decoding of PAC$(1024,512)$ code with the corresponding partially polarized channel capacity rate profile over BI-AWGN at $E_{b}/N_{0}=2.5~{}$dB with GA rate profile.
@@
Figure 13: Comparison of the partial optimum average path metric corresponding to the correctly decoded codewords of fast SC decoding of PAC$(128,64)$ code with the corresponding partially polarized channel capacity rate profile over BI-AWGN at $E_{b}/N_{0}=2.5~{}$dB with GA rate profile.
@@
Figure 14: Comparison of the partial average optimum path metric for correctly decoded codewords of fast SC decoding of a PAC$(128,64)$ code with the corresponding partially polarized channel capacity rate profile over BI-AWGN at $E_{b}/N_{0}=2.5~{}$dB, using the RM rate profile.
Furthermore, Fig. 14 shows the sample average as well as our proposed average of the metric function (obtained from the polarization of $I(W)$) for FSCL decoding of the PAC$(128,64)$ code, where the rate profile is obtained using RM code construction. In the next section, we plot the error-correction performance of the different codes studied and the average number of sorting operations required for the FSCL decoding of these codes using our proposed technique.
@@ V-B VPSCL
For a given threshold $P_{\text{th}}$, we also propose a varentropy-based list-pruning strategy for the SCL (VPSCL) decoding algorithm to reduce the number of calls to the sorting function. If the bit-metric function value $\phi(u_{i};\mathbf{y},\mathbf{u}^{i-1})$ is less than $-m_{i}$, we discard the corresponding path. With high probability (depending on $P_{\text{th}}$), this approach retains only the paths whose bit-metric values are close to their average. This method is a function of the polarized varentropy and, as a result, works for any rate profile, eliminating the need to tune the parameter $m_{T}$ as required in previous methods. It also suggests that the average number of sorting calls per decoded codeword is approximately equal to the number of bit-channels in $\mathcal{A}$ with high varentropy.
@@ V-C Numerical Results
Figure 15: Comparison of the Fano and conventional list decoding with our proposed list pruning of FSCL and VPSCL decoding algorithms for a PAC$(128,64)$ code when $m_{T}=-10$ and $P_{th}=10^{-6}$, respectively
Fig. 15 compares the error-correction performance of our proposed PFSCL with conventional SCL decoding for the PAC$(128,64)$ code using a list size of 8 and 32. The PAC code is constructed with the RM rate profile. In our proposed method, we prune branches where the corresponding metric function falls below $m_{T}=-10$. We also include the performance of our proposed VPSCL decoding algorithm with $P_{\text{th}}=10^{-6}$, corresponding to different list sizes. Additionally, the performance of the Fano decoding algorithm is presented for comparison in this figure. Table I shows the sample average number of sorting operations executed by our proposed PFSCL decoder of the PAC$(128,64)$ code, as plotted in Fig. 15. As this table shows, the average number of levels requiring pruning for a high $E_{b}/N_{0}$ value of 3.53.53.5~{}3.5dB is almost half that of a low $E_{b}/N_{0}$ value. The results for our proposed VPSCL decoding algorithm are also included in this table. As the results indicate, there is an improvement at lower $E_{b}/N_{0}$ values as well. We computed the $m_{i}$ values from the polarized varentropies $V(W_{N}^{(i)})$ obtained at $E_{b}/N_{0}=2.5~{}$dB.
TABLE I: Sample average number of sorting operations for our proposed PFSCL and VPSCL decoder algorithms for the PAC$(128,64)$ code as plotted in Fig. 15.
@@
Figure 16: Comparison of the Fano and conventional list decoding with our proposed pruning of FSCL with $m_{T}=-10$ and VPSCL decoding algorithms with $P_{th}=10^{-4}$ and $10^{-6}$ for a PAC$(1024,512)$ code all with a list size of $4$.
Fig. 16 compares the error-correction performance of our proposed PFSCL with conventional SCL decoding for the PAC$(1024,512)$ code using a list size of $4$. The code is constructed with the Gaussian approximation rate profile at $E_{b}/N_{0}=2.5~{}$dB value. In our proposed method, we prune branches where the corresponding metric function falls below $m_{T}=-10$. We also include the performance of our proposed VPSCL decoding algorithm with $P_{\text{th}}=10^{-6}$ and $P_{\text{th}}=10^{-4}$. Additionally, the performance of the Fano decoding algorithm is presented for comparison in this figure. Table II shows the sample average number of sorting operations executed by our proposed PFSCL decoder of the PAC$(1024,512)$ code, as plotted in Fig. 16. As this table shows, the average number of levels requiring sorting for a high $E_{b}/N_{0}$ value of 333~{}3dB is about $5\%$ of a conventional SCL decoding algorithm. The results for our proposed VPSCL decoding algorithm are also included in this table. As the results indicate, there is an improvement at lower $E_{b}/N_{0}$ values as well. We computed the $m_{i}$ values from the polarized varentropies $V(W_{N}^{(i)})$ obtained at $E_{b}/N_{0}=2.5~{}$dB.
TABLE II: Sample average number of sorting operations for our proposed FSCL and VPSCL decoder algorithms for the PAC$(1024,512)$ code as plotted in Fig. 16.
@@
# of sorting ($L=4$, $P_{th}=10^{-6}$) 121.78 119.03 116.89 113.84 110.52 106.61 101.83
# of sorting ($L=4$, $P_{th}=10^{-4}$) 82.86 76.24 67.38 55.25 42.11 40.74 41.49
Figure 17: Comparison of the Fano and conventional list decoding with our proposed pruning of FSCL and VPSCL decoding algorithms for a PAC$(64,32)$ code when $m_{T}=-10$ and $P_{th}=10^{-6}$, respectively.
Fig. 17 compares the error-correction performance of our proposed PFSCL with conventional SCL decoding for the PAC$(64,32)$ code using a list size of 8 and $32$. The PAC code is constructed with the MC rate profile at $E_{b}/N_{0}=5~{}$dB value [32]. In our proposed method, we prune branches where the corresponding metric function falls below $m_{T}=-10$. We also include the performance of our proposed VPSCL decoding algorithm with $P_{\text{th}}=10^{-6}$, corresponding to different list sizes. Additionally, the performance of the Fano decoding algorithm is presented for comparison in this figure.
Table III shows the sample average number of sorting operations executed by our proposed PFSCL decoder of the PAC$(64,32)$ code, as plotted in Fig. 17. The results for our proposed VPSCL decoding algorithm are also included in this table. We computed the $m_{i}$ values from the polarized varentropies $V(W_{N}^{(i)})$ obtained at $E_{b}/N_{0}=2.5~{}$dB.
TABLE III: Sample average number of sorting operations for our proposed PFSCL and VPSCL decoder algorithms for the PAC$(64,32)$ code as plotted in Fig. 17.
@@
Figure 18: Comparison of the Fano and conventional list decoding with our proposed pruning of FSCL and VPSCL decoding algorithms for a PAC$(128,99)$ code when $m_{T}=-15$.
Fig. 18 compares the error-correction performance of our proposed PFSCL with conventional SCL decoding for the PAC$(128,99)$ code using a list size of $8$ and $32$. The PAC code is constructed with the RM rate profile. In our proposed method, we prune branches where the corresponding metric function falls below $m_{T}=-15$. We also include the performance of our proposed VPSCL decoding algorithm with $P_{\text{th}}=10^{-6}$, corresponding to different list sizes. Additionally, the performance of the Fano decoding algorithm is presented for comparison in this figure. Table IV shows the sample average number of sorting operations executed by our proposed PFSCL decoder of the PAC$(128,99)$ code, as plotted in Fig. 18. As this table shows, the average number of levels requiring pruning for a high $E_{b}/N_{0}$ value of 555~{}5dB is about $33\%$ of a conventional SCL decoding algorithm and the number of node visits on the polarization tree is $28$ while for the conventional SCL decoding it is $254$. The results of our proposed VPSCL decoding algorithm are also included in this table. As the results indicate, there is an improvement at lower $E_{b}/N_{0}$ values as well. We computed the $m_{i}$ values from the polarized varentropies $V(W_{N}^{(i)})$ obtained at $E_{b}/N_{0}=3.5~{}$dB.
TABLE IV: Sample average number of sorting operations for our proposed PFSCL and VPSCL decoder algorithms for the PAC$(128,99)$ code as plotted in Fig. 18.
@@
Figure 19: Comparison of our proposed VPSCL decoding algorithm with different $P_{th}$ values for a PAC$(128,64)$ code against the LC-PSCL decoding algorithm introduced in [26] for a CRC-aided Polar$(128,64)$ code, both with a list size of $8$.
@@
Fig. 19 compares the error-correction performance of our proposed VPSCL decoding algorithm for a PAC(128, 64) code with the pruning algorithm introduced in [26] for a CRC-aided Polar(128, 64) code. For VPSCL algorithm the $m_{i}$ values are computed from the polarized varentropies $V(W_{N}^{(i)})$ obtained at $E_{b}/N_{0}=2.5~{}\text{dB}$. As the figure illustrates, PAC codes demonstrate better performance compared to polar code in the short block-length regime.

### v3

@@ IV Metric Polarization Estimation
In this section, we numerically evaluate the path metrics of FSCL decoding for PAC codes across different rate profiles. For a PAC(64,32)(64,32) code operating at $E_{b}/N_{0}=2.5$~dB, $\mathbb{E}[\phi(X;Y)]=0.7944$. For the Monte Carlo (MC)-based rate profile obtained in [35], Fig. 9 shows the special nodes corresponding to this code. A conventional SCL decoder would need to visit 2N2=1262N-2=126 nodes of the polarization tree. However, FSCL significantly reduces this requirement to only $22$ node visits, resulting in more than a 5-fold decrease in the number of node visits. In this figure, we also show the corresponding average value of the polarized metric functions obtained from the polarization of the mutual information $I(W)$ at each node. The leftmost leaf node of the tree is a Rate-0 node and contains 8 LLR values, leading to 8 metric functions, all of which have an average value of $0.1891$. On the other hand, the rightmost leaf node shows an average metric function value of $1$, indicating that these are reliable channels, and the decoder is most likely to detect the correct branch for them.
@@
Figure 10: Comparison of the partial optimum average path metric corresponding to the correctly decoded codewords of fast SC decoding of PAC(64,32)(64,32) code with the corresponding partially polarized channel capacity rate profile over BI-AWGN at $E_{b}/N_{0}=2.5$~dB with MC rate profile.
@@
Fig. 10 shows the sample average path metric values for the FSCL decoding of the PAC(64,32)(64,32) code, based on $10^{4}$ samples of correctly decoded codewords. This figure also includes the cumulative values of the average of our proposed polarized metric values derived from the polarization tree shown in Fig. 9. For the cumulative values, at the $j$th level of decoding, the average of the path metric $\sum_{i=1}^{j}\phi(U_{i};\mathbf{Y},\mathbf{U}^{i-1})$ random variable is $\sum_{i=1}^{j}I(W_{N}^{(i)})$. It can be observed that the proposed partially polarized average metric provides an excellent estimation of the actual sample average metric function values. This suggests that these values can be effectively used for normalizing metric function values in various practical settings as well.
@@
Figure 11: Correct path metric sample variance corresponding to the correctly decoded codewords of fast SC decoding of PAC(64,32)(64,32) code of Fig. 10 over BI-AWGN at $E_{b}/N_{0}=2.5$~dB with MC rate profile.
@@
Fig. 11 also demonstrates the metric variance of the PAC(64,32)(64,32) from Fig. 10 for the data bits. We observe that the variance at the data bits is nearly zero, suggesting that the average value of the metric function serves as an effective approximation.
@@
Figure 12: Comparison of the partial optimum average path metric corresponding to the correctly decoded codewords of fast SC decoding of PAC(1024,512)(1024,512) code with the corresponding partially polarized channel capacity rate profile over BI-AWGN at $E_{b}/N_{0}=2.5$~dB with GA rate profile.
@@
Figure 13: Comparison of the partial optimum average path metric corresponding to the correctly decoded codewords of fast SC decoding of PAC(128,64)(128,64) code with the corresponding partially polarized channel capacity rate profile over BI-AWGN at $E_{b}/N_{0}=2.5$~dB with GA rate profile.
@@
Figure 14: Comparison of the partial average optimum path metric for correctly decoded codewords of fast SC decoding of a PAC(128,64)(128,64) code with the corresponding partially polarized channel capacity rate profile over BI-AWGN at $E_{b}/N_{0}=2.5$~dB, using the RM rate profile.
Furthermore, Fig. 14 shows the sample average as well as our proposed average of the metric function (obtained from the polarization of $I(W)$) for FSCL decoding of the PAC(128,64)(128,64) code, where the rate profile is obtained using RM code construction. In the next section, we plot the error-correction performance of the different codes studied and the average number of sorting operations required for the FSCL decoding of these codes using our proposed technique.
@@ V-B VPSCL
For a given threshold $P_{\text{th}}$, we also propose a varentropy-based list-pruning strategy for the SCL (VPSCL) decoding algorithm to reduce the number of calls to the sorting function. If the bit-metric function value $\phi(u_{i};\mathbf{y},\mathbf{u}^{i-1})$ is less than $m_{i}$, we discard the corresponding path. With high probability (depending on $P_{\text{th}}$), this approach retains only the paths whose bit-metric values are close to their average. This method is a function of the polarized varentropy and, as a result, works for any rate profile, eliminating the need to tune the parameter $m_{T}$ as required in previous methods. It also suggests that the average number of sorting calls per decoded codeword is approximately equal to the number of bit-channels in $\mathcal{A}$ with high varentropy.
@@ V-C Numerical Results
Figure 15: Comparison of the Fano and conventional list decoding with our proposed list pruning of FSCL and VPSCL decoding algorithms for a PAC(128,64)(128,64) code when $m_{T}=-10$ and $P_{th}=10^{-6}$, respectively
Fig. 15 compares the error-correction performance of our proposed PFSCL with conventional SCL decoding for the PAC(128,64)(128,64) code using a list size of 8 and 32. The PAC code is constructed with the RM rate profile. In our proposed method, we prune branches where the corresponding metric function falls below $m_{T}=-10$. PFSCL decoding requires only $38$ time steps, compared to the $254$ time steps needed by conventional SCL decoding. We also include the performance of our proposed VPSCL decoding algorithm with $P_{\text{th}}=10^{-6}$, corresponding to different list sizes. Additionally, the performance of the Fano decoding algorithm is presented for comparison in this figure. In all the numerical results presented in our paper, we use the threshold spacing value $\Delta=2$ in the Fano decoding algorithm [18, Ch. 6], [20]. Table I shows the sample average number of sorting operations executed by our proposed PFSCL decoder of the PAC(128,64)(128,64) code, as plotted in Fig. 15. As this table shows, the average number of levels requiring pruning for a high $E_{b}/N_{0}$ value of 3.53.5~dB is almost half that of a low $E_{b}/N_{0}$ value. The results for our proposed VPSCL decoding algorithm are also included in this table. As the results indicate, there is an improvement at lower $E_{b}/N_{0}$ values as well. We computed the $m_{i}$ values from the polarized varentropies $V(W_{N}^{(i)})$ obtained at $E_{b}/N_{0}=2.5$~dB.
TABLE I: Sample average number of sorting operations for our proposed PFSCL and VPSCL decoder algorithms for the PAC(128,64)(128,64) code as plotted in Fig. 15.
@@
Figure 16: Comparison of the Fano and conventional list decoding with our proposed pruning of FSCL with $m_{T}=-10$ and VPSCL decoding algorithms with $P_{th}=10^{-4}$ and $10^{-6}$ for a PAC(1024,512)(1024,512) code all with a list size of $4$.
Fig. 16 compares the error-correction performance of our proposed PFSCL with conventional SCL decoding for the PAC(1024,512)(1024,512) code using a list size of $4$. The code is constructed with the Gaussian approximation rate profile at $E_{b}/N_{0}=2.5$~dB value. In our proposed method, we prune branches where the corresponding metric function falls below $m_{T}=-10$. PFSCL decoding requires only $162$ time steps, compared to the $2046$ time steps needed by conventional SCL decoding. We also include the performance of our proposed VPSCL decoding algorithm with $P_{\text{th}}=10^{-6}$ and $P_{\text{th}}=10^{-4}$. Additionally, the performance of the Fano decoding algorithm is presented for comparison in this figure. Table II shows the sample average number of sorting operations executed by our proposed PFSCL decoder of the PAC(1024,512)(1024,512) code, as plotted in Fig. 16. As this table shows, the average number of levels requiring sorting for a high $E_{b}/N_{0}$ value of 33~dB is about $5\%$ of a conventional SCL decoding algorithm. The results for our proposed VPSCL decoding algorithm are also included in this table. As the results indicate, there is an improvement at lower $E_{b}/N_{0}$ values as well. We computed the $m_{i}$ values from the polarized varentropies $V(W_{N}^{(i)})$ obtained at $E_{b}/N_{0}=2.5$~dB.
TABLE II: Sample average number of sorting operations for our proposed FSCL and VPSCL decoder algorithms for the PAC(1024,512)(1024,512) code as plotted in Fig. 16.
@@
# of sorting ($L=4$, $P_{th}=10^{-6}$) 114.33 111.43 108.00 104.71 100.81 96.20 89.79
# of sorting ($L=4$, $P_{th}=10^{-4}$) 73.45 66.02 55.50 41.83 37.10 40.85 41.19
Figure 17: Comparison of the Fano and conventional list decoding with our proposed pruning of FSCL and VPSCL decoding algorithms for a PAC(64,32)(64,32) code when $m_{T}=-10$ and $P_{th}=10^{-6}$, respectively
Fig. 17 compares the error-correction performance of our proposed PFSCL with conventional SCL decoding for the PAC(64,32)(64,32) code using a list size of 8 and $32$. The PAC code is constructed with the MC rate profile at $E_{b}/N_{0}=5$~dB value [35]. In our proposed method, we prune branches where the corresponding metric function falls below $m_{T}=-10$. PFSCL decoding requires only $22$ time steps, compared to the $126$ time steps needed by conventional SCL decoding. We also include the performance of our proposed VPSCL decoding algorithm with $P_{\text{th}}=10^{-6}$, corresponding to different list sizes. Additionally, the performance of the Fano decoding algorithm is presented for comparison in this figure. Table III shows the sample average number of sorting operations executed by our proposed PFSCL decoder of the PAC(64,32)(64,32) code, as plotted in Fig. 17. The results for our proposed VPSCL decoding algorithm are also included in this table. We computed the $m_{i}$ values from the polarized varentropies $V(W_{N}^{(i)})$ obtained at $E_{b}/N_{0}=2.5$~dB.
TABLE III: Sample average number of sorting operations for our proposed PFSCL and VPSCL decoder algorithms for the PAC(64,32)(64,32) code as plotted in Fig. 17.
@@
Figure 18: Comparison of the Fano and conventional list decoding with our proposed pruning of FSCL decoding algorithm for a PAC(128,99)(128,99) code when $m_{T}=-15$.
Fig. 18 compares the error-correction performance of our proposed PFSCL with conventional SCL decoding for the PAC(128,99)(128,99) code using a list size of $8$ and $32$. The PAC code is constructed with the RM rate profile. In our proposed method, we prune branches where the corresponding metric function falls below $m_{T}=-15$. PFSCL decoding requires only $28$ time steps, compared to the $254$ time steps needed by conventional SCL decoding. We also include the performance of our proposed VPSCL decoding algorithm with $P_{\text{th}}=10^{-6}$, corresponding to different list sizes. Additionally, the performance of the Fano decoding algorithm is presented for comparison in this figure. Table IV shows the sample average number of sorting operations executed by our proposed PFSCL decoder of the PAC(128,99)(128,99) code, as plotted in Fig. 18. As this table shows, the average number of levels requiring pruning for a high $E_{b}/N_{0}$ value of 55~dB is about $33\%$ of a conventional SCL decoding algorithm and the number of node visits on the polarization tree is $28$ while for the conventional SCL decoding it is $254$. The results for our proposed VPSCL decoding algorithm are also included in this table. As the results indicate, there is an improvement at lower $E_{b}/N_{0}$ values as well. We computed the $m_{i}$ values from the polarized varentropies $V(W_{N}^{(i)})$ obtained at $E_{b}/N_{0}=3.5$~dB.
TABLE IV: Sample average number of sorting operations for our proposed PFSCL and VPSCL decoder algorithms for the PAC(128,99)(128,99) code as plotted in Fig. 18.
@@
Figure 19: Comparison of our proposed VPSCL decoding algorithm with different $P_{th}$ values for a PAC(128,64)(128,64) code against the LC-PSCL decoding algorithm introduced in [28] for a CRC-aided Polar(128,64)(128,64) code, both with a list size of $8$.
@@
Figure 21: Comparison of conventional stack and list decoding with our proposed pruning strategies—VPSCL and VPStack—for polar and PAC $(128,64)$ codes. For VPSCL, we use $P_{\text{th}}=10^{-6}$, and for VPStack, we use $P_{\text{th}}=10^{-4}$.
TABLE V: Sample average number of sorting operations for our proposed VPSCL decoder algorithm for the PAC and polar(128,64)(128,64) codes as plotted in Fig. 21.
$E_{b}/N_{0}$ [dB] 0.0 0.5 1.0 1.5 2.0 2.5 3.0 3.5
# of sorting (polar, $L=32$, $P_{th}=10^{-6}$) 30.87 30.54 30.13 29.61 29.12 28.57 28.09 27.68
# of sorting (polar, $L=128$, $P_{th}=10^{-6}$) 30.14 30.00 29.84 29.68 29.48 29.27 29.06 28.85
# of sorting (PAC, $L=128$, $P_{th}=10^{-6}$) 30.46 30.48 30.56 30.70 30.91 31.19 31.47 31.74
Figure 22: Comparison of the average number of paths explored by conventional stack decoding and our proposed VPSCL pruning strategy for the PAC $(128,64)$ code with $P_{\text{th}}=10^{-4}$.
Fig. 19 compares the error-correction performance of our proposed VPSCL decoding algorithm for a PAC(128, 64) code with the pruning algorithm introduced in [28] for a CRC-aided Polar(128, 64) code. For VPSCL algorithm the $m_{i}$ values are computed from the polarized varentropies $V(W_{N}^{(i)})$ obtained at $E_{b}/N_{0}=2.5~\text{dB}$. In this example, the PAC code demonstrates better performance compared to the polar code.
@@
In Fig. 21, we compare the performance of VPSCL and SCL decoding for polar and PAC codes with various list sizes, alongside the performance of VPStack and conventional stack decoding for PAC codes. As shown in the figure, PAC codes decoded with list decoding using a list size of 128 can achieve performance comparable to that of sequential decoding (stack decoding) of PAC codes. Table V also shows the sample average number sorting operations executed by our proposed VPSCL decoder of the polar and PAC $(128,64)$ codes correpsinding to Fig. 21. Although PAC codes has a much better coding gain compared to the polar codes, the average number of levels requiring sorting for both polar and PAC codes are the same indicating that this depends on the number of bit-channels with a high varentropy. Fig. 22 also compares the average number of paths in the stack for our proposed VPStack algorithm and the conventional stack decoding algorithm, corresponding to the performance results shown in Fig. 21.